The payback method helps firms establish and identify a maximum acceptable payback period that helps in capital budgeting decisions. There are two versions of the payback method: the conventional payback method and the discounted payback method.

Consider the following case:

Green Caterpillar Garden Supplies Inc. is a small firm, and several of its managers are worried about how soon the firm will be able to recover its initial investment from Project Delta’s expected future cash flows. To answer this question, Green Caterpillar’s CFO has asked that you compute the project’s payback period using the following expected net cash flows and assuming that the cash flows are received evenly throughout each year.

Complete the following table and compute the project’s conventional payback period. Round the payback period to the nearest two decimal places. Be sure to complete the entire table—even if the values exceed the point at which the cost of the project is recovered.

Year 0

Year 1

Year 2

Year 3

Expected cash flow $-6,000,000 $2,400,000 $5,100,000 $2,100,000
Cumulative cash flow
Conventional payback period: years

The conventional payback period ignores the time value of money, and this concerns Green Caterpillar’s CFO. He has now asked you to compute Delta’s discounted payback period, assuming the company has a 7% cost of capital.

Complete the following table and perform any necessary calculations. Round the discounted cash flow values to the nearest whole dollar, and the discounted payback period to the nearest two decimal places. Again, be sure to complete the entire table—even if the values exceed the point at which the cost of the project is recovered.

Year 0

Year 1

Year 2

Year 3

Cash flow $-6,000,000 $2,400,000 $5,100,000 $2,100,000
Discounted cash flow
Cumulative discounted cash flow
Discounted payback period: years

Which version of a project’s payback period should the CFO use when evaluating Project Delta, given its theoretical superiority?

The discounted payback period

The regular payback period
Answer #1

Cumulative cash flows will be as follows:

Year 0 =-$6,000,000

Year 1 = -6,000,000+2,400,000 = -$3,600,000

Year 2 =$1,500,000

Year 3 =$3,600,000

Conventional payback period = 1 + 3,600,000/5,100,000

= 1.7059 years

Discounted payback period:

Year 0 1 2 3
Cash flows -6,000,000 2,400,000 5,100,000 2,100,000
PVF 1 0.9346 0.8734 0.8163
Discounted cash flows -6,000,000 2,243,040 4,454,340 1,714,230
Cumulative discounted cash flows -6,000,000 -3,756,960 697,380 2,411,610

Discounted payback period = 1+ 3756960/4454340

= 1.8434 years

The discounted payback period as it considers time value of money